Lisaneedham:

How many intersections are there of the graphs of the equations below? One-halfx + 5y = 6 3x + 30y = 36 A. none B. one C. two D. infinitely many

darkknight:

2 linear functions can never intersect twice, only 0 times, 1 time or infinitely many (if they overlap) so eliminate the second answer choice

darkknight:

One-halfx + 5y = 6 3x + 30y = 36 1/2x + 5y = 6 3x+ 30y = 36 we can simplify the bottom equation by dividing both sides by 3 x+10y = 12 solving for y, we have 10y = -x+12 (solve for y now) For the first equation 1/2x+5y = 6 5y = -1/2x + 6 (solve for y now) Can you solve for y in the above 2 equations and see what the equations simplify down too?

Lisaneedham:

ok thanks